Evalute limits involving the by constructjng their respective table of values



1.lim t -> 0 t/(sin t)



2.lim t -> 0 (sin t)/(2t)•(1 - cot t)/t



3.lim t -> 0 (e ^ t - 1/t)




1
Expert's answer
2022-03-31T08:51:52-0400
  1. "\lim_{t\rightarrow0}\frac{t}{sin\,t}=\lim_{t\rightarrow0}(\frac{sin\,t}{t})^{-1}=1." It contains the well-known limit "\frac{sin\,t}{t}" The values of function "\frac{t}{sin\,t}" near are: "t=0.05,\frac{t}{sin\,t}\approx1.0004;" "t=0.03,\frac{t}{sin\,t}\approx1.0002;" "t=0.01;\frac{t}{sin\,t}\approx1.00002." From the latter we can conclude that the limit is 1. Another way to compute the limit is to use L'Hôpital's rule.
  2. "\lim_{t\rightarrow0}\frac{sin\,t}{2t}\frac{1- cot\,t}{t}=+\infty" . The first fraction is the well-known limit multiplied by "\frac{1}{2}". The second fraction can be rewritten in the form: "\frac{1-cot\,t}{t}=\frac{1-\frac{cos\,t}{sin\,t}}{t}=\frac{cos\,t-sin\,t}{t\,sint}". This fraction approaches infinity (the term "cos\,t-sin\,t" approaches 1 and the term "t\,sint" approaches 0). Another way to compute the limit is to use L'Hôpital's rule.
  3. "\lim_{t\rightarrow0}\frac{e^t-1}{t}=1". It is the well-known known limit "\frac{e^t-1}{t}" . The values of function "\frac{e^t-1}{t}" near are: "t=0.05", "\frac{e^t-1}{t}\approx1.0152;" "t=0.03,\frac{e^t-1}{t}\approx1.0152;" "t=0.01," "\frac{e^t-1}{t}\approx1.0050." Another way to compute the limit is to use L'Hôpital's rule.